REACTION DYNAMICS OF THE BROMINE-BROMOFORM COMPLEX IN SOLUTION

\underline{THOMAS J. PRESTON}, MAITREYA DUTTA and F. FLEMING CRIM, The University of Wisconsin-Madison, Department of Chemistry, 1101 University Avenue, Madison, WI 53706.

We have followed the evolution of the bromine species that arise from the photolysis of bromoform in the condensed phase. Solvent caging promotes the formation of iso-bromoform (CHBr_{2}-Br), which can then release a Br atom by breaking the newly formed Br-Br bond. This ejected Br can form a van der Waals complex (Br-CHBr_{3}) with a nearby un-photolyzed bromoform molecule, which is stable during our 1 ns time-window. Using the van der Waals complex as a reservoir for Br atoms, we now proceed to drive hydrogen abstraction from CHBr_{3} by Br. Estimates indicate that the barrier to this reaction is a few thousand wavenumbers. Our goal is to introduce excitation into the C-H stretching motion of a nearby solvent CHBr_{3} to access the activated complex region of the bimolecular potential energy surface.